An urn consists of 8 balls including 5 white balls and 3 black balls. 4 balls are drawn at random. What is the probability that exactly 2 balls are black?
1
Expert's answer
2019-07-25T13:58:38-0400

P = (C32 * C52)/C84 = 3*10/70 = 3/7